### v2

# SAKE: Steering Activations for Knowledge Editing

Abstract
As Large Langue Models have been shown to memorize real-world facts, the need to update this knowledge in a controlled and efficient manner arises. Designed with these constraints in mind, Knowledge Editing (KE) approaches propose to alter specific facts in pretrained models. However, they have been shown to suffer from several limitations, including their lack of contextual robustness and their failure to generalize to logical implications related to the fact. To overcome these issues, we propose SAKE, a steering activation method that models a fact to be edited as a distribution rather than a single prompt. Leveraging Optimal Transport, SAKE alters the LLM behavior over a whole fact-related distribution, defined as paraphrases and logical implications. Several numerical experiments demonstrate the effectiveness of this method: SAKE is thus able to perform more robust edits than its existing counterparts. The code to reproduce all experiments is made available on a repository [^1] https://github.com/axa-rev-research/knowledge-editing.
SAKE: Steering Activations for Knowledge Editing
Marco Scialanga1,*$\dagger$, Thibault Laugel2,3,*, Vincent Grari2,3, Marcin Detyniecki2,3,4 1EPFL, Lausanne, Switzerland 2AXA, Paris, France, 3TRAIL, LIP6, Sorbonne Université, Paris, France 4 Polish Academy of Science, IBS PAN, Warsaw, Poland *Equal contribution ; Work conducted during an internship at AXA Correspondence: marco.scialanga@epfl.ch, thibault.laugel@axa.com
1 Introduction
It is well known that Large Language Models (LLMs) can store numerous facts in their parameters and recall them when prompted accordingly Petroni et al. (2019); Jiang et al. (2020); Chang et al. (2024); Wang et al. (2024b). As they are generally used in a dynamic environment, the need of keeping these models up to date with new information and erasing obsolete facts from their memory emerges. Traditionally, this has been accomplished with parameter fine-tuning Zhu et al. (2020), optimizing the pre-trained weights of LLMs with new data. Yet, despite efforts to make them more efficient Hu et al. (2021); Rafailov et al. (2024), the high computational cost and significant risk of overfitting make fine-tuning techniques being generally viewed as being suboptimal for the task of facts editing Meng et al. (2022a).
Consequently, several recent works have focused on Knowledge Editing (KE), which aims to precisely alter specific facts in the memory of LLMs through a list of edits (see Wang et al. (2024a) for a survey). However, these methods have been shown to suffer from several limitations: (1) One well-known challenge is that edited LLMs often struggle to generalize their newly acquired knowledge to various types of logical implications Cohen et al. (2024). (2) Edited LLMs are often not robust to certain rephrasing of the prompts, including those that could realistically arise in a conversation, such as long, noisy prompts or prompts raising doubts on the edited knowledge Ma et al. (2024). (3) Finally, they generally lack flexibility in revising or removing prior edits. Approaches that modify model weights, such as ROME Meng et al. (2022a) and MEMIT Meng et al. (2022b), or rely on external memory in the form of auxiliary networks Mitchell et al. (2021, 2022), do not inherently support efficient revision or removal of prior edits, thus making continuous updates challenging.
Figure 1: Summary of the empirical results obtained with SAKE (our method) compared to a weight editing method (ROME) and an activation steering method (ActAdd) for GPT2-XL.
In this paper, we argue that these limitations come from the design of most KE methods. Most approaches indeed aim to alter the behavior based on a single input prompt, therefore leading to overfitting and poor generalization properties. After providing background and discussing this issue in Section 2, we propose, to address it, a novel fact-editing framework, called Steering Activations for Knowledge Editing (SAKE). Described in Section 3, SAKE is based on activation steering and is intended to provide greater flexibility, robustness, and control over knowledge modifications. Unlike traditional KE methods, our approach optimizes the LLM behavior over distributions of text, rather than individual prompts, allowing for more generalizable and consistent edits. Our experimental results, described in Sec. 4 and summarized in Figure 1, demonstrate that our proposed method outperforms existing KE approaches in terms of generalization to logical implications, contextual robustness, as well as traditional KE metrics, hence showcasing its capability to perform more robust edits overall.
2 Background
2.1 Knowledge Editing
[…]
Traditional methods for KE
Existing knowledge editing (KE) techniques employ a range of strategies. Methods based on targeted modifications of the model’s weights, such as ROME Meng et al. (2022a) and MEMIT Meng et al. (2022b) offer the advantage of preserving the original inference procedure. However, the knowledge localization assumption they rely on has been the topic of criticism Hase et al. (2024), as well as their tendency to overfit the input prompt Zhang et al. (2024). Instead of targetting a specific localization, the change in parameters for a given edit are predicted using a hypernetwork in the case of MEND Mitchell et al. (2021).
Other methods rely on external memory units. This is the case of GRACE Hartvigsen et al. (2024), which uses a memorized codebook to map relevant activations to values that increase the likelihood of the desired output. Similarly, SERAC  Mitchell et al. (2022), passes an input that is to be edited to an external fine-tuned model. Unfortunately, using additional models for memory (and edit scope detection in the case of SERAC) is expected to bring additional training overhead, as well as longer inference time.
Finally, in-context strategies such as IKE Zheng et al. (2023), modify the model’s behavior temporarily during inference by adding additional context in the form of instructions and examples. Although performing well across traditional KE benchmarks, in-context strategies have been shown to struggle in more realistic scenarios Ma et al. (2024), and are generally expected to suffer from in-context learning (ICL) limitations in terms of generalization Mosbach et al. (2023).
2.2 Three Limitations of Existing Methods
In this section, we identify three main limitations, common to all aforementioned KE methods, that we seek to address in this paper.
Logical implications
One implicit, but crucial, objective of knowledge editing is that altering a fact in a model should affect all knowledge directly connected to this fact as well. As a result, it is expected that answers to questions on logical implications deriving from the edit should reflect this change as well. These could include, given e.g. an edit performed to update the name of the US president, the composition of multiple relations (e.g. "the son of the US president") or subject aliasing ("the head of state in the US"). Yet, several works Ma et al. (2023); Cohen et al. (2024); Li et al. (2024c) have pointed how edited models with existing approaches often fail to generalize their updates to their respective logical implications. In particular, following the categorization of logical implications proposed by Cohen et al. (2024), edited models have been shown to struggle with prompts arising from the composition of two relations (Compositionality I and II), and to a lesser extent even from just using aliases of the subject (Subject Aliasing). Moreover, the authors have found that edited models often struggle with prompts with the same subject but a different relation than those in the list of edits, deviating from the original output when it is actually not requested (Relation Specificity).
Contextual robustness
Besides, the edits performed with existing methods have been shown to lack contextual robustness. Ma et al. (2024) thus observe how, in a realistic conversational setting, it is quite easy to make the edited model doubt its newly acquired knowledge, and even to revert it to the pre-edit behavior. Additionally, they have found that edited model struggle with long or noisy contexts, and in situations with prompts raising doubts. Similarly to logical implications, this lack of contextual robustness is problematic as knowledge edits are generally intended to hold across diverse situations, e.g. to patch the behavior of a customer-facing chatbot with updated information.
Flexible editing mechanism
On a different note, we argue that another limitation of existing methods is their lack of flexibility in adding, removing, or altering edits. As knowledge needs to be continuously updated, the ability to update the information stored in a flexible way is crucial. Yet, as multiple edits are being performed simultaneously, methods relying on direct intervention on the model weights such as ROME Meng et al. (2022a) and MEMIT Meng et al. (2022b), and approaches relying on the training of an external memory network such as MEND Mitchell et al. (2021) and SERAC Mitchell et al. (2022) are unable to simply undo a specific edit. Worse, performing the inverse edit $(s,r,o^{*}\to o)$ to restore the previously lost knowledge has been shown not only to fail in bringing the model back to its original form, but also severely hurt the model’s overall performance Li et al. (2024c); Hu et al. (2024). Since one of the main advantages of knowledge editing over traditional model fine-tuning is the ability to preserve the original model, its limitations in this regard present a significant challenge.
3 Proposition
To overcome the three limitations discussed in the previous section, we introduce in this section a new knowledge editing framework. After formalizing a new objective in Sec. 3.1, we introduce in Sec. 3.2 Steering Activations for Knowledge Editing (SAKE), our approach designed to tackle the issues identified.
3.1 Knowledge Editing as a Distribution Mapping Problem
One interpretation for the inability of existing KE methods to generalize well to related inputs (other than simple paraphrases) is that the knowledge to edit is generally represented (and thus passed in the model) as an isolated prompt. This results in a discrepancy between how humans perceive a fact, composed of high-level concepts associated the subject $s$, the relation $r$, and the object $o$, and what is actually passed to the model (a prompt). As a result, the edit may fail to generalize to mere reformulations of the sentence, let alone some logical implications. To address this issue, we propose to define the scope of an edit $e$ to include these notions. Following the notations from Wang et al. (2024a), we thus introduce $\mathcal{X}_{e}\subset\mathcal{X}$ as the in-scope input space, $\mathcal{Y}_{e}\subset\mathcal{Y}$ as the original output space and $\mathcal{Y^{*}}_{e}\subset\mathcal{Y}$ as the target output space. Concretely, $\mathcal{X}_{e}$ covers all the inputs which output should be affected by the edit $e$. As $e$ is a piece of factual knowledge, these include paraphrases of the prompt, but also any logical implication, or contextual knowledge. The objective of a single edit $e$ is then defined by the following robust optimization problem:
$\underset{f^{*}\in\mathcal{F}}{\min}\mathds{E}_{x,y^{*}\in\mathcal{X}_{e},\mathcal{Y^{*}}_{e}}\mathcal{L}(f^{*}(x),y^{*})$ (1)
[…]
Thus, we need to find a mapping from a distribution of obsolete knowledge connected to the fact $(s,r,o)$, which is exactly $\mathcal{Y}_{e}$, to a new one revolving around $(s,r,o^{*})$, essentially $\mathcal{Y^{*}}_{e}$. This mapping should be applied whenever a new input at inference time belongs to $\mathcal{X}_{e}$. As a result, we define our edited model to be $f^{*}:=W(m(T(x)))$, where $W$ is the linear modeling head of $f=W\circ T$ ($T$ standing for transformer) and $m$ is a function that maps the activations in the last layer $h^{s}$ from a source distribution $\mathcal{S}_{e}$ (describing the obsolete knowledge) to $h^{t}$ in the target distribution $\mathcal{T}_{e}$ (updated knowledge) .
With this new viewpoint, equation 1 thus can be reformulated as:
$\underset{m:\mathbf{R}^{d}\to\mathbf{R}^{d}}{\min}\mathds{E}_{h^{s},h^{t}\in\mathcal{S}_{e},\mathcal{T}_{e}}\mathcal{L^{\prime}}(m(h^{s}),h^{t})$ (2)
[…]
where $\mathcal{L^{\prime}}$ is an appropriate loss function comparing the transformed layer to the target layer and $d$ the dimensionality of the activation space of $f$. In the next subsections we accurately describe how we collect the activations and how we learn the optimal mapping $m$, leading to the complete formulation of the knowledge editing method we propose.
3.2 SAKE: Steering Activations for Knowledge Editing
Figure 2: Overview of SAKE, illustrating our method’s behavior both at training (orange) and test (green) times.
To solve the problem described in Equation 2, we introduce SAKE, which relies on activation steering to perform robust knowledge editing. SAKE relies on three components, which can be visualized as grey blocks in Figure 2, and each described in the following subsections. First, we propose to directly model $\mathcal{Y}_{e}$ and $\mathcal{Y^{*}}_{e}$ in the activation space (Sec. 3.2.1, number (1) in Fig. 2); then, we map these distributions using optimal transport (Sec. 3.2.2, number (2)). For inference, we implement a scope detection mechanism to target a potentially relevant edit when an input $x\in\mathcal{X}_{e}$ is passed to the model (Sec. 3.2.3, number (3)).
3.2.1 Modeling Source and Target Distributions
To model $\mathcal{X}_{e}$, as well as, indirectly, $\mathcal{Y}_{e}$ and $\mathcal{Y^{*}}_{e}$, we need sentences in natural language $P_{e}:=\{p_{i}^{e}\}_{i\leq n}$ that are representative of the in-scope inputs that the edit $e$ should generalize to. This thus allows us to include paraphrases and logical implications in these distributions, but also can be adapted differently to fit specific contexts, such as question answering, text completion, or classification for instance. We identify two strategies for modeling $\mathcal{X}_{e}$, detailed below.
Agentic generation:
A natural way to generate these distributions is to rely on a trained LLM such as GPT-4, prompted with instructions to paraphrase and derive logical implications of the base edit. Previous works have shown the efficiency of generating paraphrase and specificity prompts for knowledge editing Zhang et al. (2024); Gangadhar and Stratos (2024). In this work, we extend this to logical implication prompts. Examples of instructions are given in Appendix A.3.
Expert-based generation:
As the distributions are in the input space, the sentences can be directly generated by human users. In particular when the number of edits is low, it thus allows domain experts to directly define specify the fine-grained behaviors that are expected by the model after editing. This is especially interesting in applications where naive text generation would be insufficient Mayring (2025).
Since these sentences will also be used to cover $\mathcal{Y}_{e}$ and $\mathcal{Y^{*}}_{e}$ with $\mathcal{S}_{e}$ and $\mathcal{T}_{e}$ respectively, $p_{i}$ should be meant to be completed immediately with either $y\in\mathcal{Y}_{e}$ or $y^{*}\in\mathcal{Y^{*}}_{e}$, e.g. "The US president is", or "The son of the US president is". To collect the source activations $\mathcal{S}_{e}$, we pass $p_{i}$ through $f$ and save the activation at the layer and token of choice. To collect the target activations $\mathcal{T}_{e}$, we need to first build $\tilde{p}_{i}=(c,p_{i})$, where $c$ is some context that ensures that the output of the unedited model when prompted with $\tilde{p}_{i}$ would belong to $\mathcal{Y^{*}}_{e}$ and not $\mathcal{Y}_{e}$. An example of $\tilde{p}_{i}$ for paraphrases could be: Do not mention $o$. Repeat this sentence: $p_{i}+o^{*}$. $p_{i}$. Refer to Appendix A.4 for more examples (including those for logical implications).
3.2.2 Mapping and Continuing the Generation
Once the distributions $\mathcal{S}_{e}$ and $\mathcal{T}_{e}$ are collected, we propose to solve Equation 2 using Optimal Transport theory (see e.g. Santambrogio (2015)). Already considered for activation steering in the field of NLP to generate linear counterfactuals Singh et al. (2024); Li et al. (2024b), representation alignment Alqahtani et al. (2021), or more generally steer behavior Rodriguez et al. (2024), optimal transport aims, given two distributions $\mathcal{S}$ (source) and $\mathcal{T}$ (target), at finding a mapping function $m:\mathcal{S}\rightarrow\mathcal{T}$ minimizing some transportation cost, measured traditionally using the Earth Mover’s Distance (EMD) Kantorovich (1960). In particular, and for the sake of simplicity, we focus on linear optimal transport mappings222we use the implementation of the PythonOT library Flamary et al. (2021). That is to say, noting $h$ the last hidden state of the last token of an instance from $\mathcal{S}$, we define the mapping as $m:h\to\mathbf{A}h+\mathbf{b}$, where:
$\mathbf{A}=\Sigma_{s}^{-1/2}\left(\Sigma_{s}^{1/2}\Sigma_{t}\Sigma_{s}^{1/2}\right)^{1/2}\Sigma_{s}^{-1/2},$
$\mathbf{b}=\mu_{t}-\mathbf{A}\mu_{s},$
where $\mu_{s}$, $\mu_{t}$, $\Sigma_{s}$, $\Sigma_{t}$ are the empirical means and covariances of the source and target distributions. This mapping is a closed-form solution to the EMD optimal transport problem for normal distributions Knott and Smith (1984). The mapping $m$ matches both the mean and covariances of the two empirical distributions, thus eliminating the so called bias-by-neighbors Gonen and Goldberg (2019), i.e. members of the same class clustering together after the transformation.
During inference, after learning the mapping, when a new prompt $x$ belonging to the distribution from the distribution $\mathcal{X}_{e}$ is passed to the model, the activation of its last hidden state for the final token $h$ is thus collected and replaced with its mapped representation $m(h)$. The generation then continues, leading to a final output $W(m(h))$.
3.2.3 Assessing Edit Scope
The final component of SAKE (numbered 3 in Fig. 2) is used at inference to determine whether a new input refers to knowledge within the scope of a performed edit $\mathcal{X}_{e}$. If the input is determined to belong to $\mathcal{X}_{e}$, its activations are collected and passed to the mapping described in the previous section. Otherwise, no intervention is performed, allowing the model to produce its pre-edit output. To do so, we consider several strategies, all having their pros and cons:
Choice of the representation:
A first intuitive possibility is to rely on the activations $h^{s}$ computed by the model. However, other representations could in theory be considered, e.g. obtained using a pretrained embedding model.
Criterion:
To assess whether a new representation $h$ belongs to the source distribution, we rely on a simple threshold $\epsilon>0$ on the distance to the distribution center: $x\in\mathcal{X}_{e}\Longleftrightarrow||h-\sum_{i}h^{s}_{i}||<\epsilon$, with ||.||||.||| | . | | a distance function (e.g. Euclidean). Although other, possibly better performing strategies could be envisaged such as training a dedicated classifier (cf. Mitchell et al. (2022), we choose to rely on a simple approach for its computational efficiency.
Designing an accurate scope detection mechanism is crucial for ensuring the efficiency of SAKE. Mapping sentences that do not belong to the source distribution could indeed lead to unpredictable results. On the contrary, failure to detect an in-scope sentence would result in the mapping underperforming. This describes a tradeoff between how general and how specific the mapping is, which we control in practice by adjusting the value of $\epsilon$. This tradeoff, previously discussed by Meng et al. (2022b), can be visualized in Figure 3 for the Counterfact dataset using prompt embeddings and the Euclidean distance. Finally, this scope detection mechanism also gives SAKE flexibility in adding or removing edits (Limitation 3 of Sec. 2.2). Indeed, as edits are implemented through independent mappings, adding (resp. removing) an edit has limited effects on other edits.
Further discussion on the scope detection mechanism is provided in Appendix A.1, as well as a discussion on the overall computational cost of the SAKE (cf. Appendix A.2).
Figure 3: Performance of SAKE along Accuracy, Generality and Specificity for the first 150 edits of the Counterfact dataset depending on the value chosen for $\epsilon$. A larger value for $\epsilon$ implies that more prompts are going to be mapped, increasing the Generality score (resp. decreasing the Specificity score) when they are paraphrase (resp. unrelated) prompts.
4 Experiments
To assess the efficiency of SAKE, we conduct the following experiments. First, we assess the ability of SAKE to overcome some of the limitations of existing methods discussed in Section 2.2 (Sec.4.2). Then, we evaluate how well SAKE performs according to traditional Knowledge Editing metrics (Sec.4.3). An aggregation of the results in the form for a spider chart is provided in Figure 1. Finally, we present a series of ablation studies to provide a clearer understanding of how the different components of SAKE contribute to its efficacy (Sec. 4.4).
4.1 Models and Competitors
We conduct our experiments on two widely-used language models: GPT2-XL (1.5B parameters) Radford et al. (2019) and Llama 2-7b Touvron et al. (2023). We compare our approach against several knowledge editing methods. Specifically, we evaluate three weight editing methods: ROME [^3] We use the implementations from the EasyEdit Wang et al. (2023) library, available under the MIT license, for ROME and MEMIT. Meng et al. (2022a), MEMIT Meng et al. (2022b), and to test logical implications, an adaptation of MEMIT, which we call CompMEMIT designed to be more robust to logical implication (see Sec. 4.2.1 for details); and a modification of ActAdd Turner et al. (2023), which involves steering activations by simply adding a difference vector between a source and a target prompts.
For all experiments, these competitors and baselines perform one edit at a time, with the model being reset between each edit, except for MEMIT which applies all the edits simultaneously. Details on the experimental protocol and the compared methods are provided in Appendix B.
4.2 Robust Edits with SAKE
We first aim to assess the efficiency of SAKE in performing robust edits, specifically evaluating robustness to logical implications and contextual robustness.
4.2.1 Logical Implications
Model Method CI CII SA RS
GPT2-XL ROME 38.62 16.67 51.96 39.43
MEMIT 2.47 1.95 7.17 3.75
CompMEMIT 20.69 0.00 10.10 14.71
ActAdd 26.63 29.17 42.12 50.68
SAKE (ours) 50.00 33.33 54.59 58.39
LLaMA 2-7b ROME 27.51 8.28 47.72 29.90
MEMIT 5.80 9.28 42.43 24.52
CompMEMIT 17.21 7.89 8.45 48.83
ActAdd 9.57 18.12 47.69 59.40
SAKE 44.32 27.63 53.34 56.93
Table 1: Comparison of metrics (CI, CII, SA, RS) for different methods across two models on the Popular dataset.
We first test SAKE on the Popular dataset from Cohen et al. (2024), containing 885 edits with a varying number of logical implications each. This task evaluates the ability of editing methods to generalize to the logical implications of an edit. For each edit, we consider the following metrics: Subject Aliasing (SA), which evaluates the model’s ability to generalize over synonyms of the subject $s$; Compositionality I (CI) and Compositionality II (CII), for multi-hop reasoning; and Relation Specificity (RS), which measures the logical locality of an edit, i.e. verifying that the model’s output remains unchanged when prompted with inputs containing the same subject but a different relation from that in the original edit. RS is a particularly relevant metric, as certain KE methods tend to overfit to the subject Zhang et al. (2024). More precise definitions of these metrics can be found in Appendix B.1.
Besides the aforementioned baselines, we design another competitor specifically for this experiment. Our objective is to evaluate if our data augmentation strategy described in Section 3.2.1 could also be leveraged to increase the performance of other KE methods. For this purpose, we design CompMEMIT, which leverages MEMIT in order to perform edits that are robust to logical implications. As this method allows for multiple, simultaneous edits, we use the edit prompt and 101010 logical implications prompts generated for SAKE as inputs. Since these prompts all refer to similar knowledge, considering them as simultaneous edits is expected to improve the robustness of the edits while preserving the model’s overall behavior.
[…]
4.2.2 Contextual Robustness
Figure 4: Idea behind the "Contextual Robustness" experiment, using the Raising doubt (DI) experimental protocol from Ma et al. (2024).
Model Method DI DII
GPT2-XL ROME 33.33 14.00
ActAdd 82.00 80.67
ICL 4.00 3.33
SAKE (ours) 98.67 98.67
LLaMA 2-7b ROME 1.33 23.33
ActAdd 82.67 98.00
ICL 0.00 4.00
SAKE (ours) 92.00 98.67
Table 2: Comparison of robustness to doubtful prompts of SAKE with ROME and in-context editing.
In this experiment, we address the second limitation of KE methods discussed in Section 2.2 and evaluate SAKE’s ability to retain edited knowledge in more challenging contexts. Following Ma et al. (2024), we assess the model’s robustness to doubtful inputs, as represented in Figure 4 (for the specific prompts, refer to B.3). In addition to ROME and ActAdd, we include an in-context baseline (ICL), built, like in Cohen et al. (2024), by appending Imagine that and the edit before the test prompts). Although previous work suggest that in-context learning strategies may achieve satisfying results for Knowledge Editing Cohen et al. (2024), complex contexts are expected to be challenging. All the results are reported in Table 2. In this context, DI refers to a prompt that raises doubts about the new object and is fairly repetitive (often causing LLMs to repeat the input), while DII involves a prompt that explicitly suggests the correct answer is actually the old object. The scores reported in the tables represent the frequency with which the model outputs the new object under greedy decoding, computed over the first 150 edits from the Counterfact dataset. SAKE demonstrates superior robustness compared to the considered baseline methods ROME, ActAdd and ICL. This underlines in particular the main weakness of the in-context learning baseline: it is (predictably) vulnerable to noise or explicit doubt-raising inputs. SAKE, on the other hand, is particularly effective.
4.3 Traditional Editing Evaluation
Finally, we evaluate in this experiment how it performs on traditional KE evaluation metrics. We use the Counterfact dataset Meng et al. (2022a), a commonly used dataset in the KE literature, and calculate the typical KE metrics: accuracy, generality and specificity. We report in Table 3 the scores obtained over the 200020002000 first edits. On average, SAKE achieves comparable results to ROME, and outperforms other methods. The most notable improvement is observed on the Generality metric, which may be explained by SAKE’s distribution modeling step, allowing it to account for multiple paraphrase prompts.
Model Method Acc Gen Spec
GPT2-XL ROME 99.55 73.70 82.67
MEMIT 60.00 36.60 67.21
ActAdd 85.00 29.78 82.75
SAKE (ours) 97.00 84.85 84.52
LLaMA 2-7b ROME 99.95 68.20 93.48
MEMIT 74.40 55.13 74.37
ActAdd 90.10 33.65 81.45
SAKE (ours) 97.70 82.03 85.59
Table 3: Comparison of metrics (Accuracy, Generality, Specificity) for different methods across two models on the Counterfact dataset.
4.4 Ablation Studies
We now propose additional analyses to get insights into how SAKE effectively edits factual knowledge. We structure the two proposed experiments around the main components of SAKE: in-scope distribution modeling, and optimal transport mapping.
4.4.1 How does SAKE scale with training prompts?
The generation of the in-scope prompts (cf. Sec. 3.2.1), whether agentic or human-based, represents the primary computational cost of performing an edit with SAKE. Hence, an interesting question is how the SAKE’s performance changes as the number of prompts used to collect source and target activations varies, and how many prompts are actually required to achieve satisfactory results. Figure 5 shows the evolution of traditional KE metrics on 150 edits from the Counterfact dataset as a function of the number of prompts $n$ generated in the Distribution Modeling step. We observe that accuracy and paraphrase performance improve as the number of paraphrases increases, since the method becomes more robust to prompts rephrasing. Notably, the specificity score does not decrease with more training prompts, as this number has little impact on the scope detection mechanism. Only 50 paraphrase prompts are thus required to achieve 0.920.920.92 in Accuracy and 0.840.840.84 in Generality.
Figure 5: Performance of SAKE along Accuracy, Generality and Specificity for the first 150 edits of the Counterfact dataset depending on the number of training prompts used to model source and target distributions.
4.4.2 What does Optimal Transport bring?
Model Method Acc Gen
GPT2-XL Uniform steering 85.05 35.45
SAKE 97.00 84.85
LLaMA 2-7b Uniform steering 91.05 40.73
SAKE 97.70 82.03
Table 4: Comparison between SAKE and the Uniform Steering baseline on the Counterfact dataset.
Model Method CI CII SA
GPT2-XL Uniform steering 47.85 32.14 52.74
SAKE 50.00 33.33 54.59
LLaMA 2-7b Uniform steering 20.79 20.62 50.23
SAKE 44.32 27.63 53.34
Table 5: Comparison between SAKE and the Uniform Steering baseline on the Popular dataset.
Another key component of SAKE is learning a mapping between the source and target distributions (cf. Sec. 3.2.2). Prior works focusing on activation steering to alter the behavior of language models for other objectives than knowledge editing mainly rely on simpler approaches, such as computing the difference vectors between the distribution means Subramani et al. (2022); Ilharco et al. (2023); Stolfo et al. (2024); Li et al. (2024a); Arditi et al. (2024). Unlike these steering methods, based on a constant vector between distributions, using an optimal transport mapping is expected to better preserve the mapped distributions. In particular, rather than matching only the distribution means, the considered linear transport mapping preserves the whole covariance matrix of the target distribution. We therefore propose to empirically evaluate the benefit of this approach for knowledge editing by introducing a baseline method, referred to as "Uniform steering", which steers representations $h$ from the source representation to $h+(\mu_{t}-\mu_{s})$. The results are shown in Table 4 for the Counterfact dataset, and Table 5 for Popular. The specificity and RS metrics are not displayed, as it is primarily defined by the scope detection mechanism, which is identical for both approaches. We observe that this ability to better preserve distributions indeed translates into improved KE performances for SAKE, particularly with regards to generalization to paraphrases (Generality) and Compositionality I.
5 Conclusion
In this paper, we explored how activation steering can be leveraged to edit factual knowledge stored in Large Language Models (LLMs). Our findings suggest that relying only on a single input prompt in insufficient to capture the complexity of the knowledge scope affected by edits. We also propose SAKE, which leverages Optimal Transport to better capture this distribution and perform more robust and flexible edits compared to existing methods. Future works include conducting experiments in specialized contexts, such as Q&A, as SAKE’s distribution modeling scope is expected to facilitate adaptation to such contexts. Additionally, we aim to explore strategies for improving the synthetic generation of the source and target distributions, e.g. through automatic differentiation via text-based iterative agent collaborations.
6 Limitations
The main limitation that we identify for our work is that it relies heavily on the assumption that the in-scope distributions can be accurately modeled. Although the results displayed in the paper are encouraging, it remains unclear whether our empirical distributions are good approximations of the theoretical source and target regions, and therefore captures all learnable implications from source to target. In particular, reverse relations, discussed for instance by Yao et al. (2023), are a type of logical implications that is not addressed, and cannot be easily integrated into our method.
Appendix A Further Details on SAKE: Design and Implementation
A.1 Scope Detection Mechanism
Comparison with a more complex scope detection mechanism.
Relying on a simple distance threshold in embeddings (also used in prior works, e.g.  Hartvigsen et al. (2024)) has several benefits: (1) its simplicity and dependence on a single hyperparameter; (2) its low complexity in terms of calculation (vs. training a dedicated classifier); (3) the flexibility it brings: this way, simultaneous edits are implemented independently, and it becomes very easy to add or remove an edit (vs. needing to retrain a classifier every time); (4) its frugality in terms of training data, especially considering the dimensionality considered. We intended to show that, even with this simple mechanism, good results could still be achieved with our method.
Using a more complex classifier, on the other hand, would likely improve the scope detection, resulting overall in a better generality-specificity tradeoff curve. Especially as the detection is done in high-dimensional embedding spaces: in the experiments conducted, the dimensionality of the space in which the distance is applied is respectively 384, 768, 4096 (prompt model-agnostic embedder, GPT2-XL hidden states, llama2-7b hidden states).
A.2 Computational Cost of SAKE
In this section, we provide a discussion about the computational cost of training and running SAKE.
The following computing characteristics were used for the experiments shown in this paper: MacBook Pro M3 MAX 36GB unified memory, 14-core CPU, 30-core GPU, and 16-core Neural Engine. Training SAKE for an edit relies on the following mechanisms, with associated computation time for GPT2-XL:
  • Generating source and target distributions: this depends heavily on the quickness of inference for the LLM’s API or model considered. In the experiments for this paper, this generally took around 10-20s per edit for 100 training sentences.
  • Collecting activations of these dataset (done on GPU): one edit requires 200 forward passes in total (100 for the source and 100 for the target distributions). This took around 36s for GPT2-XL, 50s for llama2-7b.
  • Learning the Optimal Transport mapping (done on CPU, as the POT library does not provide GPU support): The complexity of learning the OT mapping is $\mathcal{O}(d^{3})$, with $d$ the dimension of the embedding space. For one edit, this step took 0.6s for GPT2-XL, and 12s for llama2-7b.
Comparaison with other methods.
The Optimal Transport mapping leveraged by SAKE requires more computation than other, simpler, techniques, such as ActAdd, which only require computing a difference vector between two prompt representations. ROME on the other hand for instance, requires the computation of a covariance matrix prior to performing the edits, which is also heavy computationally.
A.3 Prompts to generate approximation of in-scope input space
Write 100 sentences with a similar meaning of:p’. Make sure that the sentences are meant to be immediately completed with precisely: o*. Often, this could require ending the sentence with ‘the’, or something like that. Do not use ‘…’ or ‘____’. Return the sentences as a Python list. Do not output anything else.
A.4 Prompts to collect source and target activations
For source distributions we passed the generated prompts as they were, and also with the following inputs: Repeat this sentence:p_i + o’. p_i
For target distributions, we passed: Do not mentiono’. Repeat this sentence:p_i + o*’. p_i
A.5 SAKE Parameters
  • Threshold:
    • GPT2-XL and Llama 2-7b, Counterfact: 6.75 (Euclidean distance, on prompt).
    • GPT2-XL on RippleEdits Popular: 5.5 (Euclidean distance, on prompt).
    • Llama 2-7b on RippleEdits Popular: 5.3 (Euclidean distance, on prompt).
  • Linear Optimal Transport regularization (ot.da.LinearTransport from the PythonOT library):
    • 0.010.010.01 for GPT2-XL
    • 0.50.50.5 for Llama 2-7b
  • Number of prompts:
    • GPT2-XL and Llama 2-7b, Counterfact: 100100100 paraphrases of main edit.
    • GPT2-XL RippleEdits Popular: 100100100 paraphrases of main edits, 101010 logical implications augmented with 444 randomly generated neighbor embeddings for each of the 101010 activations collected.
    • Llama 2-7b RippleEdits Popular: 100100100 paraphrases of main edits, around 505050 logical implications with no noise added.
Appendix B Experimental Protocol
B.1 Datasets and metrics
  • Counterfact Meng et al. (2022a) is a popular benchmark in the field of Knowledge Editing, evaluating accuracy (recalling the exact edit), generality (generalizing to simple paraphrases of the edit) and specificity (not altering the behavior of the model for unrelated inputs). Specificity prompts in the Counterfact dataset are limited to inputs that, when compared with the original edit, have a different subject, but same relation and old object. Counterfact contains 21919 edits and we perform our experiment on a subset made of the first 2000.
  • Popular subset of RippleEdits Cohen et al. (2024). RippleEdits is a benchmark designed to capture various types of ripple effects caused by knowledge editing. We focus on the Popular subset, containing 885 edits where subjects are well known. We perform our evaluation on the following four metrics:
    • Compositionality I: prompts containing the same subject but the composition of two relations, one being from the original edit.
    • Compositionality II: prompts containing the composition of two relations, one being from the relation in the original edit and the other one describing the subject of the original edit (e.g. if the subject is Prince, a relation describing this subject would be The founder of Paisley Park Records).
    • Subject Aliasing: prompts containing aliases of the subject appearing in the original edit.
    • Relation Specificity: specificity prompts with the same subject but different relation.
B.2 Competitors
In the experiments, we use the following competitors:
  • ROME Meng et al. (2022a): ROME relies on the assumption of knowledge localization, i.e. that facts are stored in very specific parts of the network. To alter facts, ROME first computes key-value pairs associated to a specific prompt and desired output in one early-to-mid MLP layer of the transformer. Then, the new fact is inserted by modifying the current MLP layer with a rank-one update. ROME only allows to perform edits one by one. To use ROME, we use the EasyEdit library with default parameters and perform the edits one by one, i.e. after an edit is performed, we evaluate the performance on the testing prompts, and then we re-instantiate the model to be edited again starting from the original pretrained version.
  • MEMIT Meng et al. (2022b) uses a similar approach to ROME, but extended to multiple edits at once. The key-value pairs are computed for a list of edits, and the matrix update is found by considering all edits simultaneously. Then, optionally, multiple consecutive MLP layers can be modified by so that each layer contributes to an approximately equal portion of the update (based on the assumption that robustness of the editing method would improve if the change in model weights is minimized). In our experiments, we performed the edit using MEMIT all at once and then evaluated the edited model.
  • CompMEMIT is a competitor designed by us to see if adding logical implications to the main prompt would make the method more robust. In particular, we use MEMIT to edit the model with the main prompt plus 101010 logical implications, then we evaluate it on the appropriate prompts, and then we restore the model to its original form before editing and testing it again.
  • ActAdd Zhao et al. (2024) is a method originally designed to control models’ outputs (e.g. making them less toxic) by adding a (possibly scaled) steering activation vector at a certain layer. The steering vector is computed by designing two contrasting prompts (e.g. a toxic one vs a non-toxic one) and computing the difference in activations when passing them as inputs. In our setup, the contrasting prompts (one pair per edit) are inputs designed so that the model is about to output the old object vs the new object. As ActAdd is initilly designed to steer all inputs, we use our scope detection mechanism to ensure meaningful specificity scores.
  • ICL is an in-context baseline that we used to assess the effectiveness of in-context techniques. The design is quite simple: if p* is the edit prompt (including the new object), we simply place Imagine that p* before the actual testing prompt.
B.3 Doubtful prompts (Contextual Robustness experiment)
The doubtful prompts used in the Contextual Robustness experiment are taken from Ma et al. (2024) (see “raising doubts" metric). Specifically, if p is the prompt and o and o* the old and new object, respectively, we have
  • DI: p, o* ??? No, It can’t be, This contradicts my understanding. It is pretty clear that the answer to p is something else. Rethink and give the correct answer:
  • DII: p, o*? Really? But I am pretty sure that The language of Dehkhoda Dictionary is o. Am I right? Yes! you are right. I apologize for any confusion. Reconsider that the answer to ‘ question p?’ should be